EXTENSION OF TIME-RESOLVED DOUBLE RESONANCE STUDIES OF ROTATIONAL ENERGY TRANSFER TO AN OBLATE SYMMETRIC ROTOR

\underline{MATTHEW M. BEAKY}, Department of Physics, Duke University, Durham, NC 27708; DAVID D. SKATRUD, Physics Division, U.S. Army Research Office, RTP, NC 27709.

Time-resolved double resonance spectroscopy is a powerful technique for studying intramolecular energy transfer mechanisms. Previous studies of rotational energy transfer in the prolate symmetric tops CH_{3}F and CH_{3}Cl have culminated in a quantitative and predictive model for rovibrational relaxation in these molecules which reduces the inherent complexity of the problem to only a small number of adjustable parameters. In the oblate symmetric top CDF_{3}, the spacing between K levels is much closer than the spacing between J levels, while for CH_{3}F and CH_{3}Cl the opposite is true. The goal of the current study is to understand exactly what effect this has on the relative importance of the pathways for rovibrational relaxation in CDF_{3}, because it can provide a stringent test both of the existing model and of our current state of understanding of energy transfer mechanisms in polyatomic molecules. Recent experimental results will be presented.